The accounting records of Whispering Winds Corp., a real estate developer, indicated income before income tax of $856,000 for its year ended December 31, 2020, and of $560,000 for the year ended December 31, 2021. The following data are also available.


Whispering Winds Corp. follows IFRS.

Part 1

Determine the balance of any deferred tax asset or liability accounts at December 31, 2019, 2020, and 2021. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Answer #1

For year 2021

a) Total reversal amount = 35100 + 36300 + 44300 = 115,700

Taxable income =Pre-tax financial income -Future taxable amount

                     = 249400-115700

                      = 133,700

Income taxes payable =Taxable income *tax rate

                      = 133,700*30%

                    = 40,110

b)

Date

Account title

Debit

credit

December 31 ,2020

Income tax expense (249400*30%)

74,820



Deferred tax liability


34,710


Income tax payable


40,110

#Out of deferred tax liability, 35,100*30% = 10,530 will be classified as current (as same is payable within one year from reporting period) and rest will be classified as non-current:34,710-10530= 24,180

c)

Income before income tax

249,400

less:Income tax expense

(74,820)

Net income

174,580

For year 2022

a) Total reversal amount = 35100 + 36300 + 44300 = 115,700

Taxable income =Pre-tax financial income -Future taxable amount

                     = 209,400-115700

                      = 93,700

Income taxes payable =Taxable income *tax rate

                      = 93,700*30%

                    = 28,110

b)

Date

Account title

Debit

credit

December 31 ,2020

Income tax expense (209400*30%)

62,820



Deferred tax liability


34,710


Income tax payable


28,110

#Out of deferred tax liability, 36,300*30% = 10,890 will be classified as current (as same is payable within one year from reporting period) and rest will be classified as non-current:34,710-10,890= 23,820

c)

Income before income tax

209,400

less: Income tax expense

(62,820)

Net income

146,580


answered by: ANURANJAN SARSAM